OpenStudy (anonymous):

7(x -2) + 3(x + 2) = 5(x - 6)

OpenStudy (anonymous):

x=-22/5

OpenStudy (jadeishere):

7x - 14 + 3x + 6 = 5x - 30 10x -8 = 5x - 30 5x = -22 x = -22/5 x = -4.4 didn't we already go over this one? o.O